# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ 2.1 Nonparametric Bounds
Now, since some conditional expectations remain latent (as shown in (2.1)), one may bound them using either known outcome supports $L^{(k)},U^{(k)}]$ or their quantile-based versions $Q_{Y}(p),Q_{Y}(1-p)]$. Manski, (1990, 2003) propose the following nonparametric bounds for the treatment effect:
$\begin{split}\Re\in\bigg{[}&\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}=1]\Pr(Z_{ijt}=1)+L^{(1)}\Pr(Z_{ijt}=0)\\ &\quad-U^{(0)}\Pr(Z_{ijt}=1)-\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}=0]\Pr(Z_{ijt}=0),\\ &\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}=1]\Pr(Z_{ijt}=1)+U^{(1)}\Pr(Z_{ijt}=0)\\ &\quad-L^{(0)}\Pr(Z_{ijt}=1)-\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}=0]\Pr(Z_{ijt}=0)\bigg{]}\end{split}$ (14)
These bounds can be tightened by substituting $L^{(k)}$ and $U^{(k)}$ with the quantiles $Q_{Y}(p)$ and $Q_{Y}(1-p)$, respectively. In essence, using a similar notation to Manski, (2003), the region $H[\Re]$ is the identification region for $\Re$, where $H[\Re]$ is defined as the bound in (14). Note that $H[\Re]$ is only partially identified when $0<\Pr[Z_{ijt}=k]<1$ for $k=0,1$, as otherwise, $H[\Re]$ is simply a singleton. In other words, if, say, $\Pr[Z_{ijt}=1]=1$, then both upper and lower bounds coincide with the treated mean, so $H[\Re]$ collapses.
In the following sections, we outline estimation procedures for both the naïve unconditional difference and the nonparametric bounds. We also construct $(1-\alpha)\%$ confidence intervals for the bounds using Bonferroni-adjusted intervals as proposed by Horowitz and Manski, (1998), and derive standard errors via the delta method [see Casella and Berger, (2024)].
@@ 2.2 Interpretation of the Bounding Constants
The bounding constants
$L^{(1)}\leq\mathbb{E}\left[Y_{ijt}^{(1)}\mid Z_{ijt}=0\right]\leq U^{(1)},\quad L^{(0)}\leq\mathbb{E}\left[Y_{ijt}^{(0)}\mid Z_{ijt}=1\right]\leq U^{(0)},$
state that the latent (never-observed) mean outcome a “treated” firm would have realised had it not been treated cannot be lower than $L^{(1)}$ nor higher than $U^{(1)}$; analogously for an “untreated” firm under treatment. Without bounding these counterfactual means the ATE, $\Re$, is not point-identifiable: any value between -$\infty$ and $\infty$ could be rationalised by suitable (and untestable) choices of $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}=0]$ and $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}=1]$. Because our outcome of interest (Tobin’s $Q$) is unbounded in theory, we adopt quantile-based limits— e.g. $L^{(k)}=Q_{Y}^{(k)}(0.10)$ and $U^{(k)}=Q_{Y}^{(k)}(0.90)$— as a reasonable compromise: they confine the worst-case counterfactual means to the central 80% of the empirical outcome distribution, ruling out only the most extreme tail behaviour while introducing minimal additional assumptions. Under these mild restrictions the interval in (14) remains robust to selection on unobservables yet is now finite, so if the entire interval lies above (below) 0 we may still conclude a positive (negative) causal effect even when ignorability fails. We therefore describe $H[\Re]$ as a set of “worst-case bounds” for the ATE under no unverifiable assumptions beyond the outcome range.
@@ 2.3 Testing in the Presence of a Random Tipping Point
As is evident from Eq. (1)–(3), the composition of treated and untreated firms depends on the threshold $\tau$. While one could fix $\tau$ and analyse the resulting samples, the goal here is different: we seek the tipping point at which the average treatment effect $\Re$ becomes strictly positive (or negative). Hence $\tau$ must be regarded as a random stopping time.
@@
Let $\mathcal{D}_{ijt}$ denote the diversity signal for firm $i$ in sector $j$ at time $t$ (for instance, the percentage of women or non-white executives). A firm is labelled “treated” when $\mathcal{D}_{ijt}\geq\tau$. Rather than prespecify $\tau$, we examine a grid of meaningful cut-offs,
$\tau_{m}=m,\quad m\in\mathcal{M}$
@@
where in our context $\mathcal{M}=\{5,10,15,\dots,90,95\}$, with $\overline{\mathcal{M}}=\lvert\mathcal{M}\rvert$ and $m_{0}=5$ and $m_{1}=95$, and, for each $m$, set
$Z_{ijt}(m)=\mathbbm{1}\{\mathcal{D}_{ijt}\geq\tau_{m}\}.$
Thus, we test
$H_{0}:0\in H[\Re_{u}]\quad\forall u\in\mathcal{M}\quad\text{against}\quad H_{1}:\exists u\in\mathcal{M}\text{ s.t. }0\notin H[\Re_{u}].$ (15)
For every $\tau_{m}$ we estimate the corresponding treatment effect using the methods in Sections 3 and 4. The selected threshold $\widehat{\tau}$ is the smallest $\tau_{m}$ whose estimated effect differs significantly from zero at the chosen level. Boundary values $\tau_{0}$ and $\tau_{100}$ are excluded, because they would collapse the identification region $H[\Re]$ to a singleton.
Following Siegmund, (2013), define the stopping rule
$\tilde{\tau}=\inf\{\tau:\tau\geq\tau_{m_{0}},(H_{*}[\Re]>0)\cup(H^{*}[\Re]<0)\}\quad(m_{0}\geq 5),$
where $H_{*}[\Re]$ and $H^{*}[\Re]$ are the lower and upper bounds of $H[\Re]$. The procedure stops at $\min(\tilde{\tau},\tau_{m_{1}})$ and rejects $H_{0}$ if $\tilde{\tau}<\tau_{m_{1}}$ and either $H_{*}[\Re]>0$ (positive effect) or $H^{*}[\Re]<0$ (negative effect).
Let $S_{N}:=(H_{*}[\Re]>0)\cup(H^{*}[\Re]<0)$. For a fixed threshold $\tau$ the test is sized so that $\Pr[S_{N}\mid H_{0}]\leq\alpha$. For the random threshold the required family-wise error bound is
$\begin{split}\Pr&\left[\left\{(\tilde{\tau}=\tau_{m_{0}})\cap S_{N}\right\}\bigcup\right.\\ \qquad&\left.\phantom{\bigcup}\left\{(\tilde{\tau}\in(\tau_{m_{0}},\tau_{m_{1}}])\cap S_{N}\right\}\bigcup\right.\\ \qquad&\left.\phantom{\bigcup}\left\{(\tilde{\tau}>\tau_{m_{1}})\cap S_{N}\right\}\bigg{|}H_{0}\right]\leq\alpha\end{split}$ (16)
which is equivalent to
$\Pr\left[\bigcup\limits_{u=m_{0}}^{m_{1}}\left\{\left(\tilde{\tau}=\tau_{u}\right)\cap S_{N}\right\}\bigcup\left\{(\tilde{\tau}>\tau_{m_{1}})\cap S_{N}\right\}\bigg{|}H_{0}\right]\leq\alpha,$ (17)
and, by Boole’s inequality,
$\begin{split}\Pr&\left[\bigcup\limits_{u=m_{0}}^{m_{1}}\left\{\left(\tilde{\tau}=\tau_{u}\right)\cap S_{N}\right\}\bigcup\left\{(\tilde{\tau}>\tau_{m_{1}})\cap S_{N}\right\}\bigg{|}H_{0}\right]\leq\\ &\qquad\phantom{\bigcup\bigcup}\sum\limits_{u=m_{0}}^{m_{1}}\Pr\left[\left(\tilde{\tau}=\tau_{u}\right)\cap S_{N}\mid H_{0}\right]+\Pr[(\tilde{\tau}>\tau_{m_{1}})\cap S_{N}\mid H_{0}]\leq\alpha,\end{split}$ (18)
so that $\sum_{u=m_{0}}^{m_{1}}\alpha_{u}+\alpha_{m_{1}+}\leq\alpha.$
Because the grid stops at its largest value $\tau_{m_{1}}$, the event $(\tilde{\tau}>\tau_{m_{1}})\cap S_{N}$ has probability zero, so the “tail” error share $\alpha_{m_{1}+}$ can be set to 0 and omitted.
@@
Since all $\overline{\mathcal{M}}=19$ looks are pre–scheduled and use the same sample, we adopt the equal–spending rule of Pocock, (1977):
[…]
@@
The two–sided stage-wise critical value is therefore $c_{u}=\Phi^{-1}\left(1-\alpha_{u}/2\right)\simeq 3.07$ when $\alpha=0.05$ and $\overline{\mathcal{M}}=19$. (Alternative allocations include O’Brien–Fleming (O’Brien and Fleming,, 1979) or the Lan–DeMets spending function (Gordon Lan and DeMets,, 1983).)

### v2

@@ 2.1 Nonparametric Bounds
Now, since some conditional expectations remain latent (as shown in (2.1)), one may bound them using either known outcome supports $L^{(k)},\;U^{(k)}]$ or their quantile-based versions $Q_{Y}(p),Q_{Y}(1-p)]$. manski1990nonparametric; manski2003partial propose the following nonparametric bounds for the treatment effect:
$\begin{split}\Re\in\bigg[&\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau)=1]\Pr(Z_{ijt}(\tau)=1)+L^{(1)}\Pr(Z_{ijt}(\tau)=0)\\ &\quad-U^{(0)}\Pr(Z_{ijt}(\tau)=1)-\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau)=0]\Pr(Z_{ijt}(\tau)=0),\\ &\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau)=1]\Pr(Z_{ijt}(\tau)=1)+U^{(1)}\Pr(Z_{ijt}(\tau)=0)\\ &\quad-L^{(0)}\Pr(Z_{ijt}(\tau)=1)-\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau)=0]\Pr(Z_{ijt}(\tau)=0)\bigg]\end{split}$ (14)
These bounds can be tightened by substituting $L^{(k)}$ and $U^{(k)}$ with the quantiles $Q_{Y}(p)$ and $Q_{Y}(1-p)$, respectively. In essence, using a similar notation to manski2003partial, the region $\mathcal{H}[\Re]$ is the identification region for $\Re$, where $\mathcal{H}[\Re]$ is defined as the bound in (14). Note that $\mathcal{H}[\Re]$ is only partially identified when $0<\Pr[Z_{ijt}(\tau)=k]<1$ for $k=0,1$, as otherwise, $\mathcal{H}[\Re]$ is simply a singleton. In other words, if, say, $\Pr[Z_{ijt}(\tau)=1]=1$, then both upper and lower bounds coincide with the treated mean, so $\mathcal{H}[\Re]$ collapses.
In the following sections, we outline estimation procedures for both the naïve unconditional difference and the nonparametric bounds. We also construct $100(1-\alpha)\%$ confidence intervals for the bounds using Bonferroni-adjusted intervals as proposed by horowitz1998censoring, and derive standard errors via the delta method [see casella2024statistical].
@@ 2.2 Interpretation of the Bounding Constants
For a fixed $\tau$, the bounding constants
$L^{(1)}\leq\mathbb{E}\left[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau)=0\right]\leq U^{(1)},\quad L^{(0)}\leq\mathbb{E}\left[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau)=1\right]\leq U^{(0)},$
state that the latent (never-observed) mean outcome a “treated” firm would have realised had it not been treated cannot be lower than $L^{(1)}$ nor higher than $U^{(1)}$; analogously for an “untreated” firm under treatment. Without bounding these counterfactual means the ATE, $\Re$, is not point-identifiable: any value between -$\infty$ and $\infty$ could be rationalised by suitable (and untestable) choices of $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau)=0]$ and $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau)=1]$. Because our outcome of interest (Tobin’s $Q$) is in theory unbounded from above, we adopt quantile-based limit, e.g. $L^{(k)}=0$ and $U^{(k)}=Q_{Y}^{(k)}(0.90)$, as a reasonable compromise: they confine the worst-case counterfactual means to the central 90% of the empirical outcome distribution, ruling out only the most extreme tail behaviour while introducing minimal additional assumptions. The 10% tails trimmed was deemed a reasonable balance between realism and conservatism in our firm performance data, which can be heavy-tailed. In Section 6, when applying Manski bounds to study the causal impact of workforce gender diversity on Tobin’s $Q$, we further experiment with tighter bounds (such as the 10th–90th percentiles) to assess how the results are affected.
Under these mild restrictions the interval in (14) remains robust to selection on unobservables yet is now finite, so if the entire interval lies above (below) 0 we may still conclude a positive (negative) causal effect even when ignorability fails. We therefore describe $\mathcal{H}[\Re]$ as a set of “worst-case bounds” for the ATE under no unverifiable assumptions beyond the outcome range.
@@ 2.3 Testing in the Presence of a Random Tipping Point
As is evident from Eq. (1)-(3), the composition of treated and untreated firms depends on the threshold $\tau$. While one could fix $\tau$ and analyse the resulting samples, our goal is different: we seek the tipping point at which the average treatment effect $\Re$ becomes strictly positive (or negative). Thus $\tau$ must be regarded as a random stopping time.
@@
Let $\mathcal{D}_{ijt}$ denote the diversity signal for firm $i$ in sector $j$ at time $t$ (e.g. the percentage of women or non-white executives). A firm is labelled “treated” when $\mathcal{D}_{ijt}\geq\tau$. Rather than prespecify $\tau$, we examine a grid of meaningful cut-offs,
$\tau_{m}=m,\qquad m\in\mathcal{M},$
@@
where in our context $\mathcal{M}=\{5,10,15,\dots,90,95\}$, and $\overline{\mathcal{M}}:=\lvert\mathcal{M}\rvert$. For each $m$ we define
$Z_{ijt}(\tau_{m})=\mathbbm{1}\{\mathcal{D}_{ijt}\geq\tau_{m}\}.$
For a chosen significance level $\alpha\in(0,1)$ we test
$H_{0}:\;0\in\mathcal{H}[\Re_{u}]\quad\forall u\in\mathcal{M}\qquad\text{vs.}\qquad H_{1}:\;\exists u\in\mathcal{M}\text{ s.t.\ }0\notin\mathcal{H}[\Re_{u}].$ (15)
Following siegmund2013sequential, the stopping rule is
$\tilde{\tau}:=\min\bigl\{\tau_{u}:\ \mathcal{H}_{*}[\Re_{u}]>0\ \text{or}\ \mathcal{H}^{*}[\Re_{u}]<0\,\bigr\},$
and we reject $H_{0}$ if $\tilde{\tau}\leq\tau_{m_{1}}$. Denote the rejection event at look $u$ by
$S^{\tau_{u}}:=\{\mathcal{H}_{*}[\Re_{u}]>0\}\cup\{\mathcal{H}^{*}[\Re_{u}]<0\},$
where $\mathcal{H}_{*}[\Re_{u}]$ and $\{\mathcal{H}^{*}[\Re_{u}]$ are the lower and upper bounds of $\mathcal{H}[\Re_{u}]$ respectively. For a fixed threshold $\tau_{u}$, the test can be sized at level $\alpha_{u}$, i.e.
$\Pr\left(S^{\tau_{u}}\mid H_{0}\right)\leq\alpha_{u}.$
For the sequential procedure, the family-wise type I error requirement is
$\Pr\left(\exists u\in\{m_{0},\dots,m_{1}\}:\ (\tilde{\tau}=\tau_{u})\cap S^{\tau_{u}}\ \Big|\ H_{0}\right)\leq\alpha.$
Since the rejection event is a union over looks,
$\{\exists u:\tilde{\tau}=\tau_{u}\text{ and }S^{\tau_{u}}\}=\bigcup_{u=m_{0}}^{m_{1}}\left\{(\tilde{\tau}=\tau_{u})\cap S^{\tau_{u}}\right\},$
Boole’s inequality yields
$\Pr\left(\bigcup_{u=m_{0}}^{m_{1}}\left\{(\tilde{\tau}=\tau_{u})\cap S^{\tau_{u}}\right\}\ \Big|\ H_{0}\right)\leq\sum_{u=m_{0}}^{m_{1}}\Pr\left((\tilde{\tau}=\tau_{u})\cap S^{\tau_{u}}\mid H_{0}\right)\leq\sum_{u=m_{0}}^{m_{1}}\alpha_{u}.$
Thus a sufficient condition for family-wise control at level $\alpha$ is
$\sum_{u=m_{0}}^{m_{1}}\alpha_{u}=\alpha.$
@@
Since all $\overline{\mathcal{M}}=19$ looks are pre-scheduled, we adopt the equal-spending rule of pocock1977group,
$\alpha_{u}=\frac{\alpha}{\overline{\mathcal{M}}},\qquad u=m_{0},\dots,m_{1}.$
@@
The corresponding two-sided critical value is $c_{u}=\Phi^{-1}(1-\alpha_{u}/2)\approx 3.007$ for $\alpha=0.05$ and $\overline{\mathcal{M}}=19$ [^1] Alternative allocations include O’Brien-Fleming (o1979multiple) or the Lan-DeMets spending function (gordon1983discrete)..
2.4 Empirical-Population CDF Divergence and Calibrated Tail Endpoints
As before, for a fixed $\tau$, let $k\in\{0,1\}$ index the untreated and treated groups respectively. For $i,j,t$ denote the population and empirical distribution functions by
$F^{(k)}(y)=\Pr\{Y_{ijt}^{(k)}\leq y\},\quad\hat{F}^{(k)}_{N_{k}}(y)=\frac{1}{N_{k}}\sum_{i,j,t}^{N_{k}}\mathbbm{1}\{Y_{ijt}^{(k)}\leq y\},$
and define the uniform deviation $\Delta_{k}:=\sup_{y\in\mathbb{R}}\lvert\hat{F}^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert$. As in Sections 2.1-2.3, the problem consists of finding suitable bounding thresholds for the latent conditional expectations $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau)=1]$ and $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau)=0]$. The problem of controlling the family-wise size $\alpha$ for a given level $\alpha\in(0,1)$ using the Bonferroni approach is discussed extensively in Section 4. However, in this Section, we discuss how to control size $\alpha^{\prime}$ for the uniform divergence of the true CDF $F^{(k)}(y)$ and its empirical counterpart $\hat{F}_{N_{k}}^{(k)}(y)$.
For group $k$, the Dvoretzky-Kiefer-Wolfowitz inequality (Massart-sharp) (dvoretzky1956asymptotic) implies that for every $t>0$,
$\Pr\left(\Delta_{k}>t\right)\leq 2\exp(-2N_{k}t^{2}).$ (16)
Selecting $t=\varepsilon_{k}$ with
$\varepsilon_{k}:=\sqrt{\frac{\log\left(2/\alpha^{\prime}\right)}{2N_{k}}}$ (17)
yields the DKW event $\mathcal{E}_{k}:=\{\Delta_{k}\leq\varepsilon_{k}\}$ with $\Pr(\mathcal{E}_{k})\geq 1-\alpha^{\prime}$. Thus,
$\Pr\left(\Delta_{k}\leq\varepsilon_{k}\right)\geq 1-\alpha^{\prime},$ (18)
so the construction is jointly valid at the targeted family-wise level.
Uniform control of the CDFs translates into control of tail quantiles. On the event $\{\Delta_{k}\leq\varepsilon_{k}\}$, for every $p\in[\varepsilon_{k},1-\varepsilon_{k}]$ and noting Lemma 1 in the Appendix, one has
$F^{-1}(p-\varepsilon_{k})\leq\hat{F}^{-1}_{N_{k}}(p)\leq F^{-1}(p+\varepsilon_{k}).$ (19)
where the superscript $(k)$ is dropped from the inverse CDFs hereafter for the ease of exposition. Choosing $p=2\varepsilon_{k}$ and $p=1-2\varepsilon_{k}$ (which lie strictly inside the admissible interval) and writing $Y^{(k)}_{(1)}\leq\cdots\leq Y^{(k)}_{(N_{k})}$ for the order statistics,
$F^{-1}(\varepsilon_{k})\leq\hat{F}^{-1}_{N_{k}}(2\varepsilon_{k})=Y^{(k)}_{(\lceil 2\varepsilon_{k}N_{k}\rceil)},\quad\hat{F}^{-1}_{N_{k}}(1-2\varepsilon_{k})=Y^{(k)}_{(\lceil(1-2\varepsilon_{k})N_{k}\rceil)}\leq F^{-1}(1-\varepsilon_{k}).$
We therefore define the DKW-calibrated tail endpoints
$L_{k}:=Y^{(k)}_{(\lceil 2\varepsilon_{k}N_{k}\rceil)},\quad U_{k}:=Y^{(k)}_{(\lceil(1-2\varepsilon_{k})N_{k}\rceil)}.$ (20)
These are “outward” in the sense that up to $\varepsilon_{k}$ probability mass is allowed to lie beyond each observed tail; the widening occurs on the probability axis via (19), and the resulting endpoints are empirical quantiles (order statistics), not additive shifts in outcome units. When population bounds $L^{(k)},U^{(k)}$ are unknown in Sections 2.1 and 2.2, we plug in $L^{(k)}\leftarrow L_{k}$ and $U^{(k)}\leftarrow U_{k}$ for the relevant group $k$.
Figure 2.1 illustrates the construction for a sample of size $N_{k}=200$. Taking $\alpha^{\prime}=0.05$ gives
$\varepsilon_{k}=\sqrt{\frac{\ln(2/0.05)}{2\cdot 200}}=0.096,$
and the shaded band $F^{(k)}\pm\varepsilon_{k}$ envelopes the empirical curve uniformly. In this case the indices in (20) are $\lceil 2\varepsilon_{k}N_{k}\rceil=39$ and $\lceil(1-2\varepsilon_{k})N_{k}\rceil=162$, which are the data-driven tail endpoints replacing unknown support extremes in the bounds and tests that follow.
Figure 2.1: Empirical and population CDFs with a DKW band for $N_{k}=200$. The shaded region is $F(y)\pm\varepsilon$ with $\varepsilon=\sqrt{\ln(2/\alpha^{\prime})/(2N_{k})}=0.096$ for $\alpha^{\prime}=0.05$.
Remark 1.
If $\varepsilon_{k}\leq\frac{1}{4}$, the convenient choice
$L_{k}=Y^{(k)}_{\left(\lceil 2\varepsilon_{k}N_{k}\rceil\right)},\quad U_{k}=Y^{(k)}_{\left(\lceil(1-2\varepsilon_{k})N_{k}\rceil\right)}$
ensures $L_{k}\leq U_{k}$. When $\varepsilon_{k}>\frac{1}{4}$, pick any tail mass level $r\in\left(0,\frac{1}{2}-\varepsilon_{k}\right]$ and set
$L_{k}=F^{-1}_{N_{k}}(r+\varepsilon_{k})=Y^{(k)}_{\left(\lceil(r+\varepsilon_{k})N_{k}\rceil\right)},\quad U_{k}=F^{-1}_{N_{k}}(1-r-\varepsilon_{k})=Y^{(k)}_{\left(\lceil(1-r-\varepsilon_{k})N_{k}\rceil\right)}.$
Then $L_{k}\leq U_{k}$ by construction, and on the DKW event $\{\sup_{y}\lvert\hat{F}^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert\leq\varepsilon_{k}\}$ we have
$F^{-1}(r)\leq L_{k}\leq U_{k}\leq F^{-1}(1-r).$
For a concrete choice that avoids ties in small samples, one may take $r=\frac{1}{2}-\varepsilon_{k}-\frac{1}{N_{k}}$ (or any $o(1)$ slack). When indices fall outside $\{1,\dots,N_{k}\}$ due to rounding, truncate them to the boundary.